the cashier at the fast food outlet rings up a bill of $24.60 for an order of 6 milkshakes and 8 burgers. the next order is for 3 milkshakes and 5 burgers and the bill is $14.25. determine the cost of each itmem.
Answer by ewatrrr(10682)

 
Hi
Let x and y represent the cost of the milkshakes and burger respectively:
Question states***
6x + 8y = 24.60
3x + 5y = 14.25
6x + 8y = 24.60
-6x - 10y = -28.50 ||Multiplying thru by -2 to eliminate x
-2y = -3.90
y = $1.95, cost of the burger
6x + 8·1.95 = 24.60
6x = 9
x = 9/6 = 3/2 = $1.50
CHECKING our Answer***
3·1.50 + 5·1.95 = $14.25